OpenStudy (anonymous):

solve the inequality x^2+x-6>or equal to-4 by using a table and a graph

OpenStudy (mr.math):

\[x^2+x-6\ge -4 \implies x^2+x-2\ge0 \implies (x+2)(x-1)\ge0 \]

OpenStudy (anonymous):

\( \Rightarrow x \le -2 \text{ or } x \gt 1 \)

OpenStudy (mr.math):

Now, you need to study the intervals as shown below: -2 1 --------------------------------- - + + x+2 - - + x-1 + - + (x+2)(x-1)

OpenStudy (mr.math):

So our intervals are then \((-\infty,-2] + [1,\infty) \).

OpenStudy (anonymous):

Just a minute typo: \[ \Rightarrow x \le -2 \text{ or } x \ge 1 \]

OpenStudy (mr.math):

Yeah.

OpenStudy (anonymous):

Thanks!

OpenStudy (mr.math):

What's the latex for union? @FFM

OpenStudy (anonymous):

\( \cup \cap \)

OpenStudy (mr.math):

lol

OpenStudy (anonymous):

;)